# M3-CVC: Controllable Video Compression with Multimodal Generative Models

### v1

@@ I Introduction
In recent years, motivated by the need for efficient compression in bandwidth-constrained environments and emerging video applications, traditional video coding standards like H.264 [1], H.265 [2] and H.266 [3] have been developed. However, these standards are built on rigid assumptions about content structure, resulting in constrained adaptability when confronted with increasingly complex video content. In this context, deep learning-based video codecs have been proposed to dynamically learn data-driven representations, allowing for enhanced signal restoration capabilities.
Deep learning-based video codecs can be broadly categorized into two types: general-purpose codecs and domain-specific codecs. General-purpose codecs are designed to compress various types of videos. Common Techniques employed include spatiotemporal contextual modeling [4, 5] and optimization utilizing generative models such as Generative Adversarial Networks (GAN) [6]. General-purpose codecs exhibit considerable versatility, while they frequently exhibit elevated compression bitrates and operate as black-box systems, rendering their internal mechanisms opaque and challenging to interpret or control directly. Domain-specific codecs focus on compressing particular types of videos, such as human face [7] and body videos [8]. Typical frameworks characterize video frames into symbolic representations including learned 2D keypoints [9] and 2D landmarks [10]. Domain-specific codecs typically achieve higher efficiency in low-bitrate compression while offering interpretable features. Nonetheless, their applicability is limited. Currently, there is a notable absence of neural video codecs possessing both general applicability and controllability while achieving extremely low bitrates. To address this gap, it is imperative to enhance the extraction of spatiotemporal and semantic information from videos and leverage this information for efficient video reconstruction [11]. The emergence of large multimodal models [12] and conditional diffusion models [13] presents significant opportunities for advancing these tasks, as they offer improved capabilities for integrating and utilizing complex information across modalities with strong priors.
@@ II-C Keyframe Compression
$\begin{split}k^{*}(u,v)&=\arg\min_{k\in\{1,\dots,l_{v}\}}d(\mathbf{h_{i}}(u,v),\mathbf{v_{k}})\\ \mathbf{K^{*}}&=\left(k^{*}(i,j)\right)_{i=1,\dots,h_{p}}^{j=1,\dots,w_{p}}\end{split}$
@@ III-A Experimental Setting
Datasets We utilize the MSR-VTT dataset [27] as the training dataset. All training videos are cropped and resized to $512\times 320$, with frame rate adjusted to 8 fps. Subsequently, several keyframes are extracted from each training video using method based on semantic difference (setting $\lambda_{2}=0$ in II-A). Pretrained LMM Qwen-VL-Instruct-2B [14] is utilized to extract visual information, ultimately producing a set of text-image pairs for end-to-end training of the image codec. As for test dataset, we select UVG [28], MCL-JCV [29], HEVC Class B and C to evaluate the performance of traditional codecs, neural codecs, and our proposed framework. All test videos are preprocessed with the same procedure as applied to the training videos.
@@
III-B Experimental Results
To comprehensively evaluate the performance of the proposed M3-CVC framework, we conduct comparisons with both traditional video codecs (H.264, H.265, and H.266) and learned methods (DCVC-DC and DCVC-FM). For H.264 and H.265, the x264 and x265 encoders are used with the ”veryslow” setting, while for H.266, we utilize VTM-17.0 under the low-delay P configuration with QP values of 63, 58, 53, 48, 43. Based on the above settings, we conducted R-D performance experiments, and the resulting R-D curves and visual results are shown in 4 and 5 respectively.. It is evident that M3-CVC outperforms the state-of-the-art VVC standard, achieving substantial bitrate savings. Notably, M3-CVC excels under extremely low bitrates, where its performance significantly surpasses that of traditional codecs. Furthermore, in terms of semantic fidelity, M3-CVC demonstrates particularly strong results, ensuring higher preservation of video content meaning while maintaining competitive performance compared to both DCVC-DC and DCVC-FM.
@@ III-C Ablation Study
To validate the effectiveness of the various methods employed in our framework, we conduct a series of ablation experiments using VTM-17.0 as the anchor. For keyframe selection strategy, we perform three comparative tests: fixed-interval keyframe extraction, semantic threshold-only and joint semantic-motion threshold. The results demonstrate that, with an equal number of keyframes, the method where both kinds of thresholds are non-zero achieves superior R-D performance. Regarding image information extraction, we compare two approaches: having the LMM describe all information in a single step versus employing a multi-turn dialogue strategy. The results confirm the effectiveness of the dialogue strategy. For video generation modes, we compare interpolation mode and prediction mode, showing that interpolation mode, which incorporates additional visual signal information, yields significantly better performance.
@@ IV Conclusion
In this paper, we introduced M3-CVC, a framework designed for general-purpose ultra-low-bitrate video compression. M3-CVC realizes controllable video information extraction by leveraging large multimodal models with specifically-designed dialogue strategies, while achieves high-fidelity video reconstruction with pretrained conditional diffusion models designed for image and video generation. Experimental results demonstrate that M3-CVC shows significant performance improvements compared to state-of-the-art technologies such as VVC, particularly under extremely low bitrate conditions.

### v2

@@ I Introduction
In recent years, motivated by the need for efficient compression in bandwidth-constrained environments and emerging video applications, traditional video coding standards like H.264(AVC)[1], H.265(HEVC)[2] and H.266(VVC)[3] have been developed. However, these standards are built on rigid assumptions about content structure, resulting in constrained adaptability when confronted with increasingly complex video content[4]. In this context, deep learning-based video codecs have been proposed to dynamically learn data-driven representations, allowing for enhanced signal restoration capabilities.
Deep learning-based video codecs can be broadly categorized into two types: general-purpose learned codecs and domain-specific codecs. General-purpose learned codecs are designed to compress various types of videos. Common Techniques employed include spatiotemporal contextual modeling[5, 6] and optimization utilizing generative models such as Generative Adversarial Networks (GAN)[7]. General-purpose learned codecs exhibit considerable versatility, while they frequently exhibit elevated compression bitrates and operate as black-box systems, rendering their internal mechanisms opaque and challenging to interpret or control directly. Domain-specific codecs focus on compressing particular types of videos[8], such as human face[9] and body videos[10]. Typical frameworks characterize video frames into symbolic representations including learned 2D keypoints[11] and 2D landmarks[12]. Domain-specific codecs typically achieve higher efficiency in low-bitrate compression while offering interpretable features. Nonetheless, their applicability is limited. Currently, there are few neural video codecs[13] possessing both general applicability and controllability while achieving extremely low bitrates. To address this gap, it is imperative to enhance the extraction of spatiotemporal and semantic information from videos and leverage this information for efficient video reconstruction[14]. The emergence of large multimodal models[15] and conditional diffusion models[16] presents significant opportunities for advancing these tasks, as they offer improved capabilities for integrating and utilizing complex information across modalities with strong priors.
@@ II-C Keyframe Compression
$\begin{split}k^{*}(u,v)&=\underset{k\in\{1,\dots,l_{v}\}}{\arg\min}\,d(\mathbf{h_{i}}(u,v),\mathbf{v_{k}})\\ \mathbf{K^{*}}&=\left(k^{*}(i,j)\right)_{i=1,\dots,h_{p}}^{j=1,\dots,w_{p}}\end{split}$
@@ III-A Experimental Setup
Datasets We utilize the MSR-VTT dataset [30] as the training dataset. All training videos are cropped and resized to $512\times 320$, with frame rate adjusted to 8 fps. Subsequently, several keyframes are extracted from each training video using method based on semantic difference (setting $\lambda_{2}=0$ in II-A). Pretrained LMM Qwen-VL-Instruct-7B [17] is utilized to extract visual information, ultimately producing a set of text-image pairs for end-to-end training of the image codec. As for test dataset, we select UVG [31], MCL-JCV [32], HEVC Class B and C to evaluate the performance of traditional codecs, learned codecs, and our proposed framework. All test videos are preprocessed with the same procedure as applied to the training videos.
@@
III-B Quantitative and Qualitative Analysis
To comprehensively evaluate the performance of the proposed M3-CVC framework, we conduct comparisons with both traditional codecs (H.264, H.265, and H.266) and learned codecs (DCVC-DC and DCVC-FM). For H.264 and H.265, FFmpeg x264 and x265 encoders are used with ”veryslow” setting. For H.266, we utilize VTM-17.0 under low-delay P configuration with QP values of 63, 58, 53, 48, 43, 37, 32, 27. Based on the above settings, we conduct R-D performance experiments, and the resulting R-D curves and visual results are shown in 4 and 5 respectively. It is evident that M3-CVC outperforms the state-of-the-art VVC standard, achieving substantial bitrate savings. Notably, M3-CVC excels under extremely low bitrates, where its performance significantly surpasses that of traditional codecs. Furthermore, in terms of semantic fidelity, M3-CVC demonstrates particularly strong results, ensuring higher preservation of video content meaning while maintaining competitive performance compared to both DCVC-DC and DCVC-FM.
@@ III-C Ablation Study
To validate the effectiveness of the methods employed in our framework, we conduct a series of ablation experiments using VTM-17.0 as the anchor. These experiments are performed on three video sequences from the HEVC Class B dataset (Kimono1, ParkScene, and Cactus). We evaluate the R-D performance using LPIPS and compute the BD-rate savings for each sequence. The overall BD-rate saving results are obtained by averaging the individual savings across all test sequences.
For keyframe selection strategy, we perform three comparative tests: fixed-interval keyframe extraction, semantic threshold-only and joint semantic-motion threshold. The results demonstrate that with an equal number of keyframes, the method where both kinds of thresholds are non-zero achieves superior R-D performance. Regarding image information extraction, we compare two approaches: having the LMM describe all information in a single step versus employing a multi-turn dialogue strategy. The results confirm the advantage of our dialogue strategy. For video generation modes, we compare interpolation mode and prediction mode, showing that interpolation mode, which incorporates additional visual signal information, yields significantly better performance.
@@
III-D Latency Performance Evaluation
We test the processing latency of M3-CVC on a single NVIDIA RTX 3090 GPU. Meanwhile, we also measure the processing latency of VTM-17.0 under low-delay P configuration on Intel Xeon Gold 6230 CPU with 40 cores as a comparison. The results are presented in Table II. The experiments are conducted on ParkScene video sequence from HEVC Class B dataset.
TABLE II: Latency Comparison of M3-CVC and VTM-17.0
Item Latency / s Item Latency / s
M3-CVC 219.2 VTM-17.0 1542.8
Encoder 76.7 Encoder 1519.7
Decoder 142.5 Decoder 23.1
The results demonstrate that M3-CVC outperforms VTM-17.0 in overall latency, particularly during encoding, due to GPU-accelerated generative models. While decoding latency is higher in M3-CVC, a common issue with diffusion-based codecs [13], the method shows promise for competitive video compression, especially when encoding latency is prioritized.
@@ IV Conclusion
In this paper, we introduce M3-CVC, a framework designed for general-purpose ultra-low-bitrate video compression. M3-CVC realizes controllable video information extraction by leveraging large multimodal models with specifically-designed dialogue strategies, while achieves high-fidelity video reconstruction with pretrained conditional diffusion models designed for image and video generation. Experimental results demonstrate that M3-CVC shows significant performance improvements compared to state-of-the-art technologies such as VVC, particularly under extremely low bitrate conditions.